# FedLPA: One-shot Federated Learning with Layer-Wise Posterior Aggregation

### v3

@@ Appendix F Comparison with the previous methods
In both the domain of natural gradient optimization [48, 62, 16] and modeling output uncertainty in variational inference [63], using the Fisher approximation of the Hessian does not involve the issue of inverting covariance. However, in the context of federated learning, when performing variational inference on the aggregated model, the necessity of inverting covariance becomes unavoidable. To address this problem, we propose a novel algorithm that constructs a quadratic objective function. During aggregation, this algorithm directly trains the aggregated model using local covariances and expectations, thereby circumventing the need for inversion operations. Additionally, we provide a detailed convergence analysis of this method in Appendix J.
Note that the previous methods [17, 51] adopt the same core approach that utilizes the online Laplace approximation to obtain diagonal Fisher for model aggregation, in which they conduct experiments on different datasets and published on different venues. We mainly analyze our approach with the comparison of DiagonalFisher [17]. DiagonalFisher assumes independence among parameters, neglecting inter-parameter correlations, resulting in inaccurate posterior approximations. However, strong correlations exist among parameters within each layer, such as matching patterns in convolutional kernels within convolutional networks. This is a crucial factor that cannot be overlooked; otherwise, aggregation of the posterior would result in lower posterior regions, as compared to our method, as illustrated in Fig 6. In complex environments, employing diagonal Fisher for aggregation would prove to be entirely ineffective, whereas our method effectively leverages inter-parameter correlations at each layer, rendering it more robust. To demonstrate, we present results comparing aggregation using diagonal Fisher and our method. We have added experiments using the settings of our paper and an MLP model (784-256-64-10) on the FMNIST dataset with three random seeds for one-shot FL, the client number is 10, and the $\beta$=0.01. The results are in Table 5.
@@
Methods like K-FAC assume direct independence among data samples to utilize expectation approximation. They obtain the inverse of Fisher from individual samples and then directly compute the expectation, thereby avoiding inverse operations. However, the expectation approximation inevitably leads to biased results during model aggregation. Detailed analysis can be found in Appendix I.
To address this issue, we propose a novel algorithm that constructs a quadratic objective function. During aggregation, this algorithm directly trains the aggregated model using local covariances and expectations, eliminating the need for inversion operations. This aims to minimize aggregation biases as much as possible, and we provide a comprehensive convergence analysis of this method in Appendix J.
@@ F.1 The efficiency of FedLPA
Appendix G Block-diagonal empirical Fisher information matrix
Storing and computing Eq. 22 for a neural network is efficient. It is equivalent to calculating the expectation of the square of the gradient of the loss function with respect to the parameters $\bm{\theta}$. Meanwhile, $\mathrm{diag}(\bar{\mathbf{H}})$ is easy to inverse, $\mathrm{diag}(\bar{\mathbf{H}})^{-1}=\frac{1}{\mathrm{diag}(\bar{\mathbf{H}})}$. Therefore, such a diagonal matrix is always used as the co-variances of a Gaussian posterior, that is:
$p(\bm{\theta}|\mathcal{D})\equiv\bm{\theta}\sim\mathcal{N}(\bm{\theta}^{*},\mathrm{diag}(\tilde{\mathbf{F}}+\lambda\mathbf{I})^{-1})$ (23)
Separately, as for layers $l$, the weight $\mathrm{vec}(\mathbf{W}_{l})$ is approximated by a Gaussian,
$\mathrm{vec}(\mathbf{W}_{l})\sim\mathcal{N}(\mathrm{vec}(\mathbf{W}_{l}^{*}),\mathrm{diag}(\tilde{\mathbf{F}}_{l}+\lambda\mathbf{I})^{-1})$ (24)
This diagonal approximation has been used in many previous works [65, 66, 50, 17], but it is crude and ignores co-relations between different parameters and only considers the self-relations of parameters. As a result, regions with high posterior probability due to strong correlation are incorrectly discarded.
To efficiently capture these correlations between different parameters, earlier studies [46, 49, 54] have proposed estimating a block-diagonal Fisher information matrix $\mathbf{F}$ instead of a coordinate-wise one.
In a $L$ layers neural network, there are $L\times L$ blocks in its empirical Fisher. For layers $l_{1}$ and $l_{2}$, their empirical Fisher block is as below:
$\tilde{\mathbf{F}}_{l_{1},l_{2}}=\mathbb{E}\left[\nabla\log p(s|\mathrm{vec}(\mathbf{W}_{l_{1}}))\nabla\log p(s|\mathrm{vec}(\mathbf{W}_{l_{2}})^{\top}\right]$ (25)
In modern back-propagation neural networks, the weights are usually updated by the back-propagation algorithm. We denote $\tilde{y}=f(x,\bm{\theta})$ as the output of the network $f(\cdot,\bm{\theta})$; $\mathcal{L}(\cdot)$ as the loss function. Given a sample $s=(x,y)$, the loss is $\mathcal{L}(y,\tilde{y})$.
After computing the derivation of loss function $g_{\mathrm{loss}}=\frac{\partial\mathcal{L}(y,f(x,\bm{\theta}))}{\partial\tilde{y}}$, the gradients used to update weights are calculated by back-propagation algorithm.
Therefore, for layers $l$, the gradients are as below,
$\mathbf{G}_{l}$ $=-\nabla\log p(s|\mathbf{W}_{l})$ (26)
$=\frac{\partial\mathcal{L}(y,f(x,\bm{\theta}))}{\partial\mathrm{vec}(\mathbf{W}_{l})}$
$=\hat{\mathbf{b}}_{l}\hat{\mathbf{a}}_{l-1}^{\top}$
where $\hat{\mathbf{b}}_{l}=\mathbf{W}_{l+1}^{\top}\hat{\mathbf{b}}_{l+1}\odot\sigma_{l}^{\prime}(\mathbf{h}_{l})$ and $\hat{\mathbf{b}}_{L}=\mathbf{W}_{L}^{\top}g_{\mathrm{loss}}\odot\sigma_{L}^{\prime}(\mathbf{h}_{L-1})$. $\odot$ is coordinate-wise multiplication. All intermediate values in Eq. 26 can be computed efficiently by back-propagation algorithm [45, 46].
Applying Eq. 26 into Eq. 25, we get a simply format of $\tilde{\mathbf{F}}_{l_{1},l_{2}}$ represented by $\mathbf{G}_{l_{1}}$ and $\mathbf{G}_{l_{2}}$,
$\tilde{\mathbf{F}}_{l_{1},l_{2}}=\mathbb{E}\left[\mathrm{vec}(\mathbf{G}_{l_{1}})\mathrm{vec}(\mathbf{G}_{l_{2}})^{\top}\right]$ (27)
Using Kronecker product property, we re-write $\mathrm{vec}(\mathbf{G}_{l})$ as a Kronecker product of $\hat{a}_{l-1}$ and $\hat{b}_{l}$,
$\mathrm{vec}(\mathbf{G}_{l})=\mathrm{vec}(\hat{\mathbf{b}}_{l}\hat{\mathbf{a}}_{l-1}^{\top})=\hat{\mathbf{a}}_{l-1}\otimes\hat{\mathbf{b}}_{l}$ (28)
Therefore, using Eq. 28, we get a new format of $\tilde{\mathbf{F}}_{l_{1},l_{2}}$,
$\tilde{\mathbf{F}}_{l_{1},l_{2}}$ $=\mathbb{E}\left[(\hat{\mathbf{a}}_{l_{1}-1}\otimes\hat{\mathbf{b}}_{l_{1}})(\hat{\mathbf{a}}_{l_{2}-1}\otimes\hat{\mathbf{b}}_{l_{2}})^{\top}\right]$ (29)
$=\mathbb{E}\left[\hat{\mathbf{a}}_{l_{1}-1}\hat{\mathbf{a}}_{l_{2}-1}^{\top}\otimes\hat{\mathbf{b}}_{l_{1}}\hat{\mathbf{b}}_{l_{2}}^{\top}\right]$
However, until the above step, there is still a significant overhead to store and inverse $\tilde{\mathbf{F}}_{l_{1},l_{2}}$. In order to save computational overhead, previous work [46] takes an approximate matrix $\hat{\mathbf{F}}_{l_{1},l_{2}}$. $\hat{\mathbf{F}}_{l_{1},l_{2}}$ is a result of approximating the expectation of the Kronecker product by the Kronecker product of expectation. The approximation is as below:
$\tilde{\mathbf{F}}_{l_{1},l_{2}}$ $=\mathbb{E}\left[\hat{\mathbf{a}}_{l_{1}-1}\hat{\mathbf{a}}_{l_{2}-1}^{\top}\otimes\hat{\mathbf{b}}_{l_{1}}\hat{\mathbf{b}}_{l_{2}}^{\top}\right]\approx\hat{\mathbf{F}}_{l_{1},l_{2}}$ (30)
$\hat{\mathbf{F}}_{l_{1},l_{2}}$ $=\mathbb{E}\left[\hat{\mathbf{a}}_{l_{1}-1}\hat{\mathbf{a}}_{l_{2}-1}^{\top}\right]\otimes\mathbb{E}\left[\hat{\mathbf{b}}_{l_{1}}\hat{\mathbf{b}}_{l_{2}}^{\top}\right]$
$=\hat{\mathbf{A}}_{l_{1}-1,l_{2}-1}\otimes\hat{\mathbf{B}}_{l_{1},l_{2}}$
where $\hat{\mathbf{A}}_{l_{1}-1,l_{2}-1}=\mathbb{E}\left[\hat{\mathbf{a}}_{l_{1}-1}\hat{\mathbf{a}}_{l_{2}-1}^{\top}\right]$ and $\hat{\mathbf{B}}_{l_{1},l_{2}}=\mathbb{E}\left[\hat{\mathbf{b}}_{l_{1}}\hat{\mathbf{b}}_{l_{2}}^{\top}\right]$.
It is important to note that, however, the expectation of a Kronecker product is not equal to the Kronecker product of expectations. Such an approximation is equivalent to assuming that $\hat{\mathbf{a}}_{l-1}$ and $\hat{\mathbf{b}}_{l}$ are independent, but it is obviously not true because they are weakly related due to back-propagation. Although the approximation is biased, it could well capture the coarse structure of the empirical Fisher information matrix as demonstrated [46] and fit in the one-shot federated learning settings to obtain the information of vital importance, which may also explain the performance result for the multiple round setting in Figure 2. The visualization of the block-diagonal Fisher information matrix is shown in Appendix G.1. Meanwhile, Eq. 30 saves the significant computational overhead of inverting Kronecker Product and small sizes of $\hat{\mathbf{A}}_{l_{1}-1,l_{2}-1}$ and $\hat{\mathbf{B}}_{l_{1},l_{2}}$.
By filling all blocks into $\tilde{\mathbf{F}}$, we obtain $\tilde{\mathbf{F}}$ in a format of the Khatri-Rao product,
$\tilde{\mathbf{F}}\approx\hat{\mathbf{F}}=\left[\begin{array}[]{cccc}\hat{\mathbf{A}}_{0,0}\otimes\hat{\mathbf{B}}_{1,1}&\hat{\mathbf{A}}_{0,1}\otimes\hat{\mathbf{B}}_{1,2}&\cdots&\hat{\mathbf{A}}_{0,L-1}\otimes\hat{\mathbf{B}}_{1,L}\\ \hat{\mathbf{A}}_{1,0}\otimes\hat{\mathbf{B}}_{2,1}&\hat{\mathbf{A}}_{1,1}\otimes\hat{\mathbf{B}}_{2,2}&\cdots&\hat{\mathbf{A}}_{1,L-1}\otimes\hat{\mathbf{B}}_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ \hat{\mathbf{A}}_{L-1,0}\otimes\hat{\mathbf{B}}_{L,1}&\hat{\mathbf{A}}_{L-1,1}\otimes\hat{\mathbf{B}}_{\ell,2}&\cdots&\hat{\mathbf{A}}_{L-1,L-1}\otimes\hat{\mathbf{B}}_{L,L}\end{array}\right]$ (31)
Therefore, the Gaussian posterior is given,
$p(\bm{\theta}|\mathcal{D})\equiv\bm{\theta}\sim\mathcal{N}(\bm{\theta}^{*},(\hat{\mathbf{F}}+\lambda\mathbf{I})^{-1})$ (32)
The prior co-variances $\lambda^{-1}\mathbf{I}$ are only added into a diagonal block of $\hat{\mathbf{F}}$, but it is not convenient to be inverted. Similar to an approximation in Eq. 30, we add prior co-variances inside Kronecker product,
$\hat{\mathbf{F}}_{l,l}+\lambda\mathbf{I}$ $=\hat{\mathbf{A}}_{l-1,l-1}\otimes\hat{\mathbf{B}}_{l,l}+\lambda\mathbf{I}\otimes\mathbf{I}$ (33)
$\approx\mathbf{F}_{l,l}$
$=(\hat{\mathbf{A}}_{l-1,l-1}+\pi_{l}\sqrt{\lambda}\mathbf{I})\otimes(\hat{\mathbf{B}}_{l,l}+\frac{1}{\pi_{l}}\sqrt{\lambda}\mathbf{I})$
$=\hat{\mathbf{A}}_{l-1,l-1}\otimes\hat{\mathbf{B}}_{l,l}+\pi_{l}\sqrt{\lambda}\mathbf{I}\otimes\hat{\mathbf{B}}_{l,l}$
$\quad\quad\quad\quad\quad+\frac{1}{\pi_{l}}\sqrt{\lambda}\hat{\mathbf{A}}_{l-1,l-1}\otimes\mathbf{I}+\lambda\mathbf{I}\otimes\mathbf{I}$
where $\pi_{l}$ is a factor minimizing approximation error in $\hat{\mathbf{F}}_{l,l}$. Previous works [46, 49] give it formula as below,
$\pi_{l}=\sqrt{\frac{\left\|\hat{\mathbf{A}}_{l-1,l-1}\otimes\mathbf{I}\right\|}{\left\|\mathbf{I}\otimes\hat{\mathbf{B}}_{l,l}\right\|}}$ (34)
Previous work [46] shows that co-relations inner a layer are much more significant than others. Therefore, it is reasonable to assume that layers are independent of each other, which means the Fisher information matrix is block-diagonal. Then the posterior for layer $l$ is given [46] as below,
$\mathrm{vec}(\mathbf{W}_{l})\sim\mathcal{N}(\mathrm{vec}(\mathbf{W}^{*}_{l}),\mathbf{F}_{l,l}^{-1})$ (35)
Additionally, keeping co-relations between layers is feasible using Eq. 31. Researchers [46] have proposed a tri-diagonal approximation for the Fisher information matrix, which considers the co-relations between two consecutive layers, but the effect is not significant compared with the required extra computational resource.
Therefore, in our methodology, we apply the layer-wise Laplace approximation to obtain a block-diagonal co-variance of the posterior.
For simplicity, for the $k$-th client, we use $\mathbf{A}_{k_{l}}=\hat{\mathbf{A}}_{l-1,l-1}+\pi_{l}\sqrt{\lambda}\mathbf{I}=\hat{\mathbf{a}}_{k_{l-1}}\hat{\mathbf{a}}_{k_{l-1}}^{\top}+\pi_{l}\sqrt{\lambda}\mathbf{I}$; We use $\mathbf{B}_{k_{l}}=\hat{\mathbf{B}}_{l,l}+\frac{1}{\pi_{l}}\sqrt{\lambda}\mathbf{I}=\hat{\mathbf{b}}_{k_{l}}\hat{\mathbf{b}}_{k_{l}}^{\top}+\frac{1}{\pi_{l}}\sqrt{\lambda}\mathbf{I}$
For layer $l$ on client $k$, its empirical Fisher block is as below,
G.1 Visualization of the block-diagonal empirical Fisher information matrix
Figure 3: Visualization of block-diagonal empirical Fisher information matrix.
In Figure 3, we give an example, showing that the block-diagonal empirical Fisher information matrix could well capture the coarse structure of the full Fisher information matrix based on a neural network with three convolution layers and one fully connected layer trained on MNIST dataset. The four diagonal blocks, corresponding to four layers, can capture the most useful information of the original Fisher Information Matrix. The approximation of the block-diagonal empirical Fisher information matrix can converge fast with most information learned and works well in the one-shot scenario. Thus, our proposed FedLPA can yield good performance with this approximation.
However, it neglects inferior information in non-diagonal blocks. When we extend FedLPA to the multi-round setting, the inferior information that we omit may enrich the learning performance. While in the one-shot setting, the most useful information of the original Fisher Information Matrix matters. That may be the reason why the multi-round results of FedLPA are worse than the alternatives in Fig 2.
Appendix H Detailed discussion for the time complexity of Eq. 12
@@ G.1 Visualization of the block-diagonal empirical Fisher information matrix
Appendix I Expectation approximation (EA)
@@
Appendix J Convergence analysis of Eq. 14
We denote $\mathbf{v}=\mathrm{vec}(\bar{\mathbf{M}})$. For $\mathbf{v}\in\mathcal{V}$, we have $\|\mathbf{v}\|^{2}\leq D$ and $D$ is a positive constant. We let $g(\mathbf{v})=\frac{1}{2}\|\sum_{k}^{K}\mathrm{vec}(\mathbf{B}_{k}\bar{\mathbf{M}}\mathbf{A}_{k})-\bar{\mathbf{z}}\|_{2}^{2}$. Then we have
$g(\mathbf{v})=\frac{1}{2}\|(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})\mathrm{vec}(\bar{\mathbf{M}})-\bar{\mathbf{z}}\|_{2}^{2}=\frac{1}{2}\|(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})\mathbf{v}-\bar{\mathbf{z}}\|_{2}^{2}$ (40)
Then we have
$\nabla_{\mathbf{v}}g(\mathbf{v})=\mathbf{v}^{\top}(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})^{\top}(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})-\bar{\mathbf{z}}^{\top}(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})$ (41)
Therefore, for any two points $\mathbf{v}_{1},\mathbf{v}_{2}$, we have
$\|\nabla_{\mathbf{v}}g(\mathbf{v}_{1})-\nabla_{\mathbf{v}}g(\mathbf{v}_{2})\|\leq\|\mathbf{v}_{1}-\mathbf{v}_{2}\|\|(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})^{\top}(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})\|$ (42)
where the inequality is due to the Cauchy-Schwarz inequality. Since $\mathbf{A}_{k}$ and $\mathbf{B}_{k}$ are all constant, there exists a positive constant $L$ s.t. $\|\nabla_{\mathbf{v}}g(\mathbf{v}_{1})-\nabla_{\mathbf{v}}g(\mathbf{v}_{2})\|\leq L\|\mathbf{v}_{1}-\mathbf{v}_{2}\|\|$. Therefore $g(\mathbf{v})$ is $L$-Lipschitz smooth. Moreover, we have $\nabla^{2}_{\mathbf{v}}g(\mathbf{v})=(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})^{\top}(\sum_{k}^{K}\mathbf{A}_{k}^{\top}\otimes\mathbf{B}_{k})$, which is positive semi-definite. So $g(\mathbf{v})$ is convex w.r.t $\mathbf{v}$. Therefore, let $\delta=\frac{1}{L}$, then we have $g(\mathbf{v}^{t+1})\leq g(\mathbf{v}^{t})-\frac{1}{2L}\|\nabla g(\mathbf{v}^{t})\|^{2}$. We have
$\|\mathbf{v}^{t+1}-\mathbf{v}^{*}\|^{2}$ $=\|\mathbf{v}^{t}-\delta\nabla g(\mathbf{v}^{t})-\mathbf{v}^{*}\|^{2}$ (43)
$=\|\mathbf{v}^{t}-\mathbf{v}^{*}\|^{2}-\frac{2}{L}(\langle\nabla g(\mathbf{v}^{t})^{\top},\mathbf{v}^{t}-\mathbf{v}^{*}\rangle-\frac{1}{2L}\|\nabla g(\mathbf{v}^{t})\|^{2})$ (44)
Then we obtain
$\sum_{t=0}^{T-1}g(\mathbf{v}^{t+1})-g(\mathbf{v}^{*})$ $\leq\sum_{t=0}^{T-1}(g(\mathbf{v}^{t})-g(\mathbf{v}^{*})-\frac{1}{2L}\|\nabla g(\mathbf{v}^{t})\|^{2})$ (45)
$\leq\sum_{t=0}^{T-1}(\langle\nabla g(\mathbf{v}^{t})^{\top},\mathbf{v}^{t}-\mathbf{v}^{*}\rangle-\frac{1}{2L}\|\nabla g(\mathbf{v}^{t})\|^{2})$ (46)
$=\frac{L}{2}\sum_{t=0}^{T-1}(\|\mathbf{v}^{t}-\mathbf{v}^{*}\|^{2}-\|\mathbf{v}^{t+1}-\mathbf{v}^{*}\|^{2})$ (47)
$\leq\frac{L}{2}\|\mathbf{v}^{0}-\mathbf{v}^{*}\|^{2}$ (48)
Therefore, we obtain $g(\mathbf{v}^{t+1})-g(\mathbf{v}^{*})\leq\frac{L}{2t}\|\mathbf{v}^{0}-\mathbf{v}^{*}\|^{2}\leq\frac{LD^{2}}{2t}$, it has a linear convergent rate. Note that $\mathbf{v}=\mathrm{vec}(\bar{\mathbf{M}})$, this also implies that we have a linear convergent rate to $\bar{\mathbf{M}}^{*}$.
We choose the $l_{2}$ squared loss because it is the most common convex function used in the machine learning field and we have also proved the converge rate of FedLPA in $l_{2}$ squared loss. We simply employ the linear programming to solve the problems, the details could be viewed in our codes line 404-427 in the expeiremnts_our.py. We calculate the gradient of this squared loss function, the time complexity is linear. Other convex loss functions will not affect the implementation of FedLPA. We leave the impact of other loss functions as future work.
Appendix K Extend FedLPA to the models with enormous single layer weight parameters
@@
Appendix L Privacay discussion of FedLPA
@@
Besides, our t-SNE illustration in Fig 1 shows the classification capability on the global model, which can separate the classes. However, our figures of the t-SNE illustrations on local models in Appendix M.2 show that for the data belonging to the same class, their t-SNE illustrations are erratically distributed on different local nodes. For instance, for node 2, its training data only has 3 classes while most of the training data locates in class 5. However, it is hard for the server to infer that label distribution since the t-SNE illustration both on node 2 and other nodes also seems irregular.
@@
FedLPA is intuitively compatible with existing privacy-preserving techniques, such as differential privacy (DP) [77, 78], secure multiparty computation (SMC) [79, 80], and homomorphic encryption (HE) [81, 82, 83]. In Appendix L.1, we propose a naive DP-FedLPA with two different mechanisms to show the compatibility with differential privacy. In Appendix L.2, using iDLG attack [76], we show that our proposed FedLPA has the same privacy-preserving level as the conventional federated learning algorithms (i.e, FedAvg, FedProx, FedNova and Dense). Compared with FedAvg, we have conducted a detailed analysis from a privacy attack perspective to show that our proposed FedLAP exhibits a security level consistent with FedAvg against several types of privacy attacks, where the details are shown in Appendix L.3. Note that the main focus of FedLPA is to improve the learning performance on the one-shot FL settings, thus, we leave the integration with other privacy-preserving techniques beyond DP as an open problem.
L.1 Experiments with differential privacy

### v5

@@ Appendix B Comparison with the previous methods
In both the domain of natural gradient optimization [48, 70, 16] and modeling output uncertainty in variational inference [71], using the Fisher approximation of the Hessian does not involve the issue of inverting covariance. However, in the context of federated learning, when performing variational inference on the aggregated model, the necessity of inverting covariance becomes unavoidable. To address this problem, we propose a novel algorithm that constructs a quadratic objective function. During aggregation, this algorithm directly trains the aggregated model using local covariances and expectations, thereby circumventing the need for inversion operations.
Note that the previous methods [51, 17] adopt the same core approach that utilizes the online Laplace approximation to obtain diagonal Fisher for model aggregation, in which they conduct experiments on different datasets and published on different venues. We mainly analyze our approach with the comparison of DiagonalFisher [17]. DiagonalFisher assumes independence among parameters, neglecting inter-parameter correlations, resulting in inaccurate posterior approximations. However, strong correlations exist among parameters within each layer, such as matching patterns in convolutional kernels within convolutional networks. This is a crucial factor that cannot be overlooked; otherwise, aggregation of the posterior would result in lower posterior regions, as compared to our method. In complex environments, employing diagonal Fisher for aggregation would prove to be entirely ineffective, whereas our method effectively leverages inter-parameter correlations at each layer, rendering it more robust. To demonstrate, we present results comparing aggregation using diagonal Fisher and our method. We have added experiments using the settings of our paper and an MLP model (784-256-64-10) on the FMNIST dataset with five random seeds for one-shot FL, the client number is 10, and the $\beta$=0.01. The results are in Table 6.
@@
Methods like K-FAC assume direct independence among data samples to utilize expectation approximation. They obtain the inverse of Fisher from individual samples and then directly compute the expectation, thereby avoiding inverse operations. However, the expectation approximation inevitably leads to biased results during model aggregation. Detailed analysis can be found in Appendix D.
To address this issue, we propose a novel algorithm that constructs a quadratic objective function. During aggregation, this algorithm directly trains the aggregated model using local covariances and expectations, eliminating the need for inversion operations. This aims to minimize aggregation biases as much as possible.
@@ B.1 The efficiency of FedLPA
Appendix C Detailed discussion for the time complexity of Eq. 12
@@
Appendix D Expectation approximation (EA)
@@
Appendix E Extend FedLPA to the models with enormous single layer weight parameters
@@
Appendix F Privacay discussion of FedLPA
@@
Besides, our t-SNE illustration in Fig 1 shows the classification capability on the global model, which can separate the classes. However, our figures of the t-SNE illustrations on local models in Appendix G.1 show that for the data belonging to the same class, their t-SNE illustrations are erratically distributed on different local nodes. For instance, for node 2, its training data only has 3 classes while most of the training data locates in class 5. However, it is hard for the server to infer that label distribution since the t-SNE illustration both on node 2 and other nodes also seems irregular.
@@
FedLPA is intuitively compatible with existing privacy-preserving techniques, such as differential privacy (DP) [56, 57], secure multiparty computation (SMC) [58, 59], and homomorphic encryption (HE) [60, 61, 62]. In Appendix F.1, we propose a naive DP-FedLPA with two different mechanisms to show the compatibility with differential privacy. In Appendix F.2, using iDLG attack [82], we show that our proposed FedLPA has the same privacy-preserving level as the conventional federated learning algorithms (i.e, FedAvg, FedProx, FedNova and Dense). Compared with FedAvg, we have conducted a detailed analysis from a privacy attack perspective to show that our proposed FedLAP exhibits a security level consistent with FedAvg against several types of privacy attacks, where the details are shown in Appendix F.3. Note that the main focus of FedLPA is to improve the learning performance on the one-shot FL settings, thus, we leave the integration with other privacy-preserving techniques beyond DP as an open problem.
F.1 Experiments with differential privacy